### v6

@@ Abstract
We study how disagreement influences team performance in a dynamic game with positive production externalities. Players can hold different views about the productivity of the available production technologies. This disagreement results in different technology and effort choices – “optimistic” views induce higher effort than “skeptical” views. Views are changed when falsified by evidence. With a single technology available, optimists exert more effort early on if the team also includes skeptics. With sufficiently strong externalities, a disagreeing team produces, on average, more than any like-minded team. With multiple technologies, disagreement over which technology works best motivates everyone to exert more effort.
@@ 1 Introduction
Seminal contributions to the economic literature have warned us about the perils of preference disagreement, shown to create impasse and inefficiencies in many domains of social interaction. [^1] For instance, disagreement has been shown to impede decision-making and compromise economic outcomes in social choice (e.g. Arrow,, 1951), communication (e.g., Crawford and Sobel,, 1982), public finance and public good provision (e.g, Alesina and Tabellini,, 1990; Alesina and La Ferrara,, 2005). Yet, history suggests a link between the conflict of worldviews and greater innovation. The development of the first iPhone was, anecdotally, a story of disagreement (Merchant,, 2018; Grant,, 2021). Steve Jobs initially thought the product would only appeal to a “pocket protector” crowd, and saw the project as a dead end. A team of hard-working engineers – with Apple’s design chief Jonathan Ive on their side – had a very different opinion. They believed that the touchscreen technology would represent a paradigm shift for the industry. Job’s skepticism meant that the team needed to design a prototype so good that it would have been impossible for him not to change his mind. [^2] According to Grant, (2021): “Fadell and his engineers chipped away at the resistance by building early prototypes in secret, showing Jobs demos, and refining their designs.” Some argue that Job’s open disagreement culture was the key to the company’s success in those years (e.g., Scott,, 2017; Grant,, 2021).
@@
Beyond the iPhone anecdote, the power of disagreement has been recognized in many innovation-related contexts. Open disagreement between scientists has led to the production of more and better theories, resulting in the belief that scientific skepticism – the tendency to challenge and falsify existing theories – lies naturally at the heart of scientific progress (Kuhn,, 1962). Peers’ skepticism has typically motivated philosophers to design sophisticated arguments in favor of their worldview, to convince others to adopt it. [^3] For instance, determined to convince skeptics about the existence of God, Anselm of Canterbury designed the first onthological argument, a class of arguments that has fascinated philosophers for almost a thousand years (see https://plato.stanford.edu/entries/ontological-arguments/). Artists have often found in partners’ disagreement and competition of ideas a motivating force inspiring them to innovate and often reach success. Salmela and Oikkonen, (2022) illustrate this point using the case study of the heavy metal band Metallica and and many of their historical producers (most notably, producer Bob Rock): “Both sides [Metallica and Rock] wanted to make the best record in the world but disagreed on what it should be like and how it should be done. […] the most constructive outcomes of conflicts realized when Metallica collaborated with its partners but at the same time also competed with them. The partners first competed to find the best idea, after which the best idea was further developed together.”
@@
Outside the realm of intellectual and artistic debates, the space race was perhaps one of the starkest examples of how the competition of conflicting models (of production and society) spurred technological investment, leading to remarkable innovations in an attempt to prove the superiority of one worldview over the other (Center and Bates,, 2009). [^4] According to Brian C. Odom, Nasa Chief Historian: “In the global South, you had a lot of countries becoming independent from former colonial powers. What system would they follow? Would they follow the U.S. liberal democracy or would they follow the Soviet example of communism? Kennedy saw the race to the moon as a way to demonstrate American technological power and the benefit of one system over another.” See https://www.space.com/space-race.html). The more recent standards wars in the Tech industry – battles for market dominance between incompatible technologies – can be thought of in a similar fashion. [^5] R&D breakthroughs are recognized as typical ways of persuading the key players and winning such wars. This was the case, for instance, in the battle between the AC and DC technologies for the genetation and distribution of power. See Shapiro and Varian, (1999) for a discussion of standard wars.
@@
We provide intuitive conditions that make the upward pressure on effort prevail, and propose a simple application to team formation: an output-maximizing manager who needs to form two teams from a pool of two optimists and two skeptics might benefit from pairing together co-workers who disagree. Following a similar logic, we show that – when production externalities are sufficiently strong – adding a skeptic to a team of optimists increases output more than adding an additional optimist, even if optimists always work more than skeptics.
We build on the main insights of the single-technology case to study the implications of disagreement when alternative production technologies are available. When two technologies are similarly effective, a group of players who disagree over which technology works best is more productive than any group of individuals who share the same model – even if the latter believe that all the technologies are highly productive. The same holds if we conjecture that one technology must be better than the other, but both seem equally promising ex-ante. By working harder in the first period, each player expects to prove – through her breakthroughs – the superiority of her production technology, convincing others to adopt it. As a result, more breakthroughs are obtained early on.
Our results rely on a key feature of the disagreeing technology views: according to both the skeptic and the optimistic view, the more effort is invested in a technology, the easier it becomes to tell which view correctly describes such technology. This requires some restrictions on the alternative views considered, [^6] It does not, however, constrain the objective stochastic process mapping effort in distributions of output. In fact, many of our results hold even if both technology views are “wrong.” that we formalize relying on the concept of effort’s Blackwell informativeness (Blackwell and Girshick,, 1962). [^7] The constraint is satisfied when players model technologies along the lines of many effort-decision problems in the literature (e.g., Keller et al.,, 2005; Heidhues et al.,, 2018; Dong,, 2018; Ba,, 2022). In the conclusive sections of the paper, we discuss the implications of changing the relationship between effort and information arrival and provide examples of pairs of technology views that yield more distinctive predictions at low effort levels. We discuss how some of our main results about the optimality of disagreeing teams change in these different economic environments, and show that many of these results are restored under additional easily interpretable changes in economic primitives.
@@ 1.1 Literature Review
To reach a more structured understanding of economic implications of imperfect models of the world and their evolution, economic theorists have recently developed foundations to incorporate model misspecification (e.g., Esponda and Pouzo,, 2016; Heidhues et al.,, 2018; Rees-Jones and Taubinsky,, 2019) and “paradigm shifts” (e.g. Hong et al.,, 2007; Ortoleva,, 2012) in games and decision theoretical problems. To the current day, most of the literature has focused on long-term beliefs (e.g. Esponda and Pouzo,, 2016; Heidhues et al.,, 2018) and long-term misspecification robustness (Fudenberg and Lanzani,, 2022; Ba,, 2022), while fewer papers have addressed the incentives arising when players think strategically about each others’ model differences and change. Galperti, (2019) shows that a persuader can exploit events that are ruled out by the model of the receiver, in order to expand the scope of persuasion. Schwartzstein and Sunderam, (2021) characterize the scope of “model persuasion,” where the sender can change the receiver’s mind by proposing new models that better fit past evidence. The present paper contributes to this stream of literature by addressing a related but different question: can the persuasion incentives stemming from model differences be leveraged to increase a team’s output?
@@
Some of our ideas are related to the literature on Bayesian learning with different priors (e.g., Hart and Rinott,, 2020; Kartik et al.,, 2021), and its implications for evidence collection and disclosure (Che and Kartik,, 2009; Van den Steen,, 2010). In particular, Che and Kartik, (2009) and Van den Steen, (2010) show that disagreement with the principal can motivate an expert with aligned preferences to collect more evidence. Differently from these two papers, we analyze a setup where the persuasion technology is productive (persuasion is achieved via successful production histories), resulting in a novel connection between disagreement and productive performance. [^8] Van den Steen, (2010) also finds that disagreement between a manager and an employee tends to reduce delegation, hinder motivation, and lower satisfaction. By focusing on different strategic incentives, we provide a complementary, more optimistic, picture of the effects of disagreement on team effort and production. We provide an argument for why an output-maximizing manager could expect to benefit from hiring a disagreeing team even if she knows the characteristics of the production technology and therefore does not value information arrival per se. In addition, we show that the persuasion incentives might reduce the productive effort of skeptics, and we highlight how the average productivity gain due to disagreement depends on what the conflicting views imply for the informativeness of effort.
@@
Our application of the single-technology case to teamwork shares qualitative similarities with the literature on exponential bandits (Keller et al.,, 2005), and in particular Dong, (2018). However, our driving mechanism is substantially different and leads to different behavioral and welfare implications. [^9] Dong, (2018) finds that asymmetric information about the risky arm can increase experimentation in a team. Her results are driven by informational asymmetries and signaling instead of open disagreement and persuasion. In Dong, (2018), the player with bad news increases effort in order to pool with the player with good news. In our model, the optimistic player takes the lead, working harder bring the skeptic on board.
@@ 2 A Simple Illustration
Ann, Bob, Tom, and Lea are four engineers working in the $R\&D$ division of a tech company. One day, Ann reaches out to her three colleagues with an idea for a new product, which – she claims – could be a groundbreaking innovation and earn their company great technological and reputational advantages. Tom is convinced by Ann’s pitch. Bob and Lea, however, are skeptical. They think that the project poses structural issues and is not worth the effort.
@@
$U^{i}(y^{i},y^{-i},e^{i})=y^{i}+\beta y^{-i}-c\frac{e^{2}}{2},$
where $y^{i}$ and $y^{-i}$ are the output of players $i$ and her co-worker respectively, and $\beta>0$ implies a positive externality from production, a reduced form to capture complementarities between players’ work or motivational benefits obtaining when fellow team members get engaged in the collaboration and obtain breakthroughs. Finally, we assume that research will take place over two periods: every period, each of the two engineers chooses an effort level $e\in[0,1]$, observes the breakthroughs of both engineers (independently drawn) and receives a payoff.
@@
Now, imagine that engineers can change their mind if proven wrong: each of them will adopt the view of the co-worker if and only if they observe production outcomes possible only under the latter’s model. In the second period, players with models $\mathcal{H}$ and $\mathcal{L}$ will still pick $e^{\mathcal{H}}$ and $e^{\mathcal{L}}$ respectively. However, the FOC for Ann’s first-period effort becomes
@@
Imagine now that Bob proposes an alternative approach for the development of the innovative product. Let $x$ denote Ann’s approach, and $y$ Bob’s proposal. Bob is optimistic about $y$, but Ann is skeptical about it. Tom is optimistic about both ideas and – after a change in mind – Lea is too. The views of the four engineers can be summarized by four tuples, respectively $(\mathcal{L}_{x},\mathcal{H}_{y}),(\mathcal{H}_{x},\mathcal{L}_{y}),(\mathcal{H}_{x},\mathcal{H}_{y})$ and $(\mathcal{H}_{x},\mathcal{H}_{y})$, which we call models. In every period, each engineer will choose (i) which approach to work on; and (ii) how much research effort to exert. Between periods, she adopts her co-worker’s model if first-period returns have zero probability based on her initial model but not the co-worker’s.
@@ 3.3 Game Timeline
At the beginning of period $t=1,2$, Ann and Bob simultaneously choose their production technologies $k_{t}^{A}$ and $k_{t}^{B}$, and effort levels $e_{t}^{A}$ and $e_{t}^{B}$, respectively. Actions $\mathbf{k}_{t}=(k_{t}^{A},k_{t}^{B})$ and $\mathbf{e}_{t}=(e_{t}^{A},e_{t}^{B})$ are publicly observed. At the end of the period, output levels $y_{t}^{A}$ and $y_{t}^{B}$ – grouped in profile $\mathbf{y}_{t}$ – are drawn from $Q_{k_{t}^{A}}(\cdot|e^{A}_{t})$ and $Q_{k_{t}^{B}}(\cdot|e^{B}_{t})$ and publicly observed. Stage payoffs are realized. The game timeline is reported in figure 1.
@@ 3.4 Model Change
Denote by $\phi_{m^{i}}^{m^{-i}}(\mathbf{e},\mathbf{k}|m^{i})$ the probability of a switch from $m^{-i}$ to $m^{i}$ conditional $Q=m^{i}$ and action profile $(\mathbf{e},\mathbf{k})$. We make the following assumption.
@@ 3.5 Hypothesis Testing Equilibrium
A stage strategy $s^{i}_{t}=(s^{i}_{te},s^{i}_{tk})$ for player $i$ consists of an effort rule $s^{i}_{te}:M^{2}\to\mathcal{E}$ and a technology rule $s^{i}_{tk}:M^{2}\to K$ specifying, respectively, $i$’s effort and technology choice for each possible pair of models held by the two players at the beginning of $t$. A strategy profile $s^{i}=(s^{i}_{1},s^{i}_{2})$ consists of a stage strategy for each $t$. We use the following equilibrium notion.191919 To simplify exposition the analysis focuses on pure strategies, but allowing for mixed strategies does not change our results. Equilibrium existence never requires randomization in the second period. Randomization is never required in equilibrium in the first period if Ann and Bob start the game with the same model. At this level of generality, in specific instances, mixing in the first period might be required for equilibrium existence when players start the game with different models. In those cases, all our propositions still hold, and the inequalities of lemma 4.2 hold for each element in the supports of Ann and Bob’s mixed strategies.
@@
so that, for instance, $\hat{Y}(m^{A},m^{B},Q)>\hat{Y}(m^{A},m^{A},Q)$ means that the expected output of the “most productive” equilibrium is higher when Ann and Bob disagree than when they share Ann’s model. [^20] For the results of this paper, focusing on the most productive equilibrium is not restrictive. Note that, for each $m,m^{\prime}\in M$, $\hat{Y}(m^{\prime},m,Q)=\hat{Y}(m,m^{\prime},Q)$, by simmetry.

### v7

@@ Abstract
We study how disagreement influences team performance in a dynamic game with positive production externalities. Players can hold different views about the productivity of the available production technologies. This disagreement results in different technology and effort choices — “optimistic” views induce higher effort than “skeptical” views. Views are changed when falsified by evidence. With a single technology available, optimists exert more effort early on if the team also includes skeptics. With sufficiently strong externalities, a disagreeing team produces, on average, more than any like-minded team. With multiple technologies, disagreement over which technology works best motivates everyone to exert more effort.
@@ 1 Introduction
Seminal contributions to the economic literature have warned us about the perils of preference disagreement, shown to create impasse and inefficiencies in many domains of social interaction. [^1] For instance, disagreement has been shown to impede decision-making and compromise economic outcomes in social choice (e.g. Arrow,, 1951), communication (e.g., Crawford and Sobel,, 1982), public finance and public good provision (e.g, Alesina and Tabellini,, 1990; Alesina and La Ferrara,, 2005). Yet, history suggests a link between the conflict of worldviews and greater innovation. The development of the first iPhone was, anecdotally, a story of disagreement (Merchant,, 2018; Grant,, 2021). Steve Jobs initially thought the product would only appeal to a “pocket protector” crowd, and saw the project as a dead end. A team of hard-working engineers — with Apple’s design chief Jonathan Ive on their side — had a very different opinion. They believed that the touchscreen technology would represent a paradigm shift for the industry. Job’s skepticism meant that the team needed to design a prototype so good that it would have been impossible for him not to change his mind. [^2] According to Grant, (2021): “Fadell and his engineers chipped away at the resistance by building early prototypes in secret, showing Jobs demos, and refining their designs.” Some argue that Job’s open disagreement culture was the key to the company’s success in those years (e.g., Scott,, 2017; Grant,, 2021).
@@
Beyond the iPhone anecdote, the power of disagreement has been recognized in many innovation-related contexts. Open disagreement between scientists has led to the production of more and better theories, resulting in the belief that scientific skepticism — the tendency to challenge and falsify existing theories — lies naturally at the heart of scientific progress (Kuhn,, 1962). Peers’ skepticism has typically motivated philosophers to design sophisticated arguments in favor of their worldview, to convince others to adopt it. [^3] For instance, determined to convince skeptics about the existence of God, Anselm of Canterbury designed the first onthological argument, a class of arguments that has fascinated philosophers for almost a thousand years (see https://plato.stanford.edu/entries/ontological-arguments/). Artists have often found in partners’ disagreement and competition of ideas a motivating force inspiring them to innovate and often reach success. Salmela and Oikkonen, (2022) illustrate this point using the case study of the heavy metal band Metallica and and many of their historical producers (most notably, producer Bob Rock): “Both sides [Metallica and Rock] wanted to make the best record in the world but disagreed on what it should be like and how it should be done. […] the most constructive outcomes of conflicts realized when Metallica collaborated with its partners but at the same time also competed with them. The partners first competed to find the best idea, after which the best idea was further developed together.”
@@
Outside the realm of intellectual and artistic debates, the space race was perhaps one of the starkest examples of how the competition of conflicting models (of production and society) spurred technological investment, leading to remarkable innovations in an attempt to prove the superiority of one worldview over the other (Center and Bates,, 2009). [^4] According to Brian C. Odom, Nasa Chief Historian: “In the global South, you had a lot of countries becoming independent from former colonial powers. What system would they follow? Would they follow the U.S. liberal democracy or would they follow the Soviet example of communism? Kennedy saw the race to the moon as a way to demonstrate American technological power and the benefit of one system over another.” See https://www.space.com/space-race.html). The more recent standards wars in the Tech industry — battles for market dominance between incompatible technologies — can be thought of in a similar fashion. [^5] R&D breakthroughs are recognized as typical ways of persuading the key players and winning such wars. This was the case, for instance, in the battle between the AC and DC technologies for the genetation and distribution of power. See Shapiro and Varian, (1999) for a discussion of standard wars.
@@
We provide intuitive conditions that make the upward pressure on effort prevail, and propose a simple application to team formation: an output-maximizing manager who needs to form two teams from a pool of two optimists and two skeptics might benefit from pairing together co-workers who disagree. Following a similar logic, we show that — when production externalities are sufficiently strong — adding a skeptic to a team of optimists increases output more than adding an additional optimist, even if optimists always work more than skeptics.
We build on the main insights of the single-technology case to study the implications of disagreement when alternative production technologies are available. When two technologies are similarly effective, a group of players who disagree over which technology works best is more productive than any group of individuals who share the same model — even if the latter believe that all the technologies are highly productive. The same holds if we conjecture that one technology must be better than the other, but both seem equally promising ex-ante. By working harder in the first period, each player expects to prove — through her breakthroughs — the superiority of her production technology, convincing others to adopt it. As a result, more breakthroughs are obtained early on.
Our results rely on a key feature of the disagreeing technology views: according to both the skeptic and the optimistic view, the more effort is invested in a technology, the easier it becomes to tell which view correctly describes such technology. This requires some restrictions on the alternative views considered, [^6] It does not, however, constrain the objective stochastic process mapping effort in distributions of output. In fact, many of our results hold even if both technology views are “wrong.” that we formalize relying on the concept of effort’s Blackwell informativeness (Blackwell and Girshick,, 1962). [^7] The constraint is satisfied when players model technologies along the lines of many effort-decision problems in the literature (e.g., Keller et al.,, 2005; Heidhues et al.,, 2018; Dong,, 2018; Ba,, 2022). In the conclusive sections of the paper, we discuss the implications of changing the relationship between effort and information arrival and provide examples of pairs of technology views that yield more distinctive predictions at low effort levels. We highlight how some of our main results about the optimality of disagreeing teams change in these different economic environments, and show that many of these results are restored under additional easily interpretable changes in economic primitives.
@@ 1.1 Literature Review
To reach a more structured understanding of the economic implications of imperfect models of the world and their evolution, economic theorists have recently developed foundations to incorporate model misspecification (e.g., Esponda and Pouzo,, 2016; Heidhues et al.,, 2018; Rees-Jones and Taubinsky,, 2019) and “paradigm shifts” (e.g. Hong et al.,, 2007; Ortoleva,, 2012) in games and decision theoretical problems. To the current day, most of the literature has focused on long-term beliefs (e.g. Esponda and Pouzo,, 2016; Heidhues et al.,, 2018) and long-term misspecification robustness (Fudenberg and Lanzani,, 2022; Ba,, 2022), while fewer papers have addressed the incentives arising when players think strategically about each others’ model differences and change. Galperti, (2019) shows that a persuader can exploit events that are ruled out by the model of the receiver, to expand the scope of persuasion. Schwartzstein and Sunderam, (2021) characterize the scope of “model persuasion,” where the sender can change the receiver’s mind by proposing new models that better fit past evidence. The present paper contributes to this stream of literature by addressing a related but different question: can the persuasion incentives stemming from model differences be leveraged to increase a team’s output?
@@
Some of our ideas are related to the literature on Bayesian learning with different priors (e.g., Hart and Rinott,, 2020; Kartik et al.,, 2021), and its applications to principal-agent problems of evidence collection (e.g., Che and Kartik,, 2009; Van den Steen,, 2010) and delegation (e.g., Hirsch,, 2016). In particular, Che and Kartik, (2009) and Van den Steen, (2010) show that disagreement with the principal can motivate an expert with aligned preferences to collect more evidence. [^8] Van den Steen, (2010) also finds that disagreement between a manager and an employee tends to reduce delegation, hinder motivation, and lower satisfaction. By focusing on different strategic incentives, we provide a complementary, more optimistic, picture of the effects of disagreement on team effort and production. We complement this literature with an analysis of team formation and teamwork, where agents persuade each other through successful production histories. On the one hand, we find that an output-maximizing manager could benefit from pairing employees with different views, even if she knows the characteristics of the production technology and is not driven by information motives. On the other hand, our analysis suggests that the persuasion incentives of disagreeing co-workers might reduce the average output of skeptical team members. We also show that the implications of disagreement change drastically when team members’ output predictions differ the most at low effort levels, as well as in the presence of negative production externalities. [^9] As discussed in section 6, a shared belief that hard work obfuscates, rather than bringing to light, information about the technology promise would lower the productivity of disagreeing teams.
@@
At a high level, our paper is related to the economic literature on team diversity (e.g., Hong and Page,, 2001, 2004; Page,, 2007; Dong and Mayskaya,, 2022). We highlight how the interaction between co-workers’ strategic incentives and their perspectives on the production process can define whether team members with different views will, on average, outperform homogeneous teams. Our application of the single-technology case to teamwork shares qualitative similarities with the literature on exponential bandits (Keller et al.,, 2005, e.g.,), and in particular Dong, (2018). However, our driving mechanism is substantially different and leads to different behavioral and welfare implications. [^10] Dong, (2018) finds that asymmetric information about the risky arm can increase experimentation in a team. Her results are driven by informational asymmetries and signaling instead of open disagreement and persuasion. In Dong, (2018), the player with bad news increases effort in order to pool with the player with good news. In our model, the optimistic player takes the lead, working harder bring the skeptic on board.
@@ 2 A Simple Illustration
Ann, Bob, Tom, and Lea are four engineers working in the $R\&D$ division of a tech company. One day, Ann reaches out to her three colleagues with an idea for a new product, which — she claims — could be a groundbreaking innovation and earn their company great technological and reputational advantages. Tom is convinced by Ann’s pitch. Bob and Lea, however, are skeptical. They think that the project poses structural issues and is not worth the effort.
@@
$U^{i}(y^{i},y^{-i},e^{i})=y^{i}+\beta y^{-i}-c\frac{(e^{i})^{2}}{2},$
where $y^{i}$ and $y^{-i}$ are the output of players $i$ and her co-worker respectively, and $\beta>0$ implies a positive externality from production, a reduced form to capture complementarities between players’ work or motivational benefits obtaining when fellow team members get engaged in the collaboration and obtain breakthroughs. Finally, we assume that research will take place over two periods: every period, each of the two engineers chooses an effort level $e\in[0,1]$, observes the (independently drawn) breakthroughs of both engineers and receives a payoff.
@@
Now, imagine that engineers can change their minds if proven wrong: each of them will adopt the view of the co-worker if and only if they observe production outcomes possible only under the latter’s model. In the second period, players with models $\mathcal{H}$ and $\mathcal{L}$ will still pick $e^{\mathcal{H}}$ and $e^{\mathcal{L}}$ respectively. However, the FOC for Ann’s first-period effort becomes
@@
Imagine now that Bob proposes an alternative approach for the development of the innovative product. Let $x$ denote Ann’s approach, and $y$ Bob’s proposal. Bob is optimistic about $y$, but Ann is skeptical about it. Tom is optimistic about both ideas and — after a change in mind — Lea is too. The views of the four engineers can be summarized by four tuples, respectively $(\mathcal{L}_{x},\mathcal{H}_{y}),(\mathcal{H}_{x},\mathcal{L}_{y}),(\mathcal{H}_{x},\mathcal{H}_{y})$ and $(\mathcal{H}_{x},\mathcal{H}_{y})$, which we call models. In every period, each engineer will choose (i) which approach to work on; and (ii) how much research effort to exert. Between periods, she adopts her co-worker’s model if first-period returns have zero probability based on her initial model but not the co-worker’s.
@@ 3.3 Game Timeline
At the beginning of period $t=1,2$, Ann and Bob simultaneously choose their production technologies $k_{t}^{A}$ and $k_{t}^{B}$, and effort levels $e_{t}^{A}$ and $e_{t}^{B}$, respectively. Actions $\mathbf{k}_{t}=(k_{t}^{A},k_{t}^{B})$ and $\mathbf{e}_{t}=(e_{t}^{A},e_{t}^{B})$ are publicly observed. At the end of the period, output levels $y_{t}^{A}$ and $y_{t}^{B}$ — grouped in profile $\mathbf{y}_{t}$ — are drawn from $Q_{k_{t}^{A}}(\cdot|e^{A}_{t})$ and $Q_{k_{t}^{B}}(\cdot|e^{B}_{t})$ and publicly observed. Stage payoffs are realized. The game timeline is reported in figure 1.
@@ 3.4 Model Change
Denote by $\phi_{m^{i}}^{m^{-i}}(\mathbf{e},\mathbf{k}|m^{i})$ the probability of a switch from $m^{-i}$ to $m^{i}$ conditional on $Q=m^{i}$ and action profile $(\mathbf{e},\mathbf{k})$. We make the following assumption.
@@ 3.5 Hypothesis Testing Equilibrium
A stage strategy $s^{i}_{t}=(s^{i}_{te},s^{i}_{tk})$ for player $i$ consists of an effort rule $s^{i}_{te}:M^{2}\to\mathcal{E}$ and a technology rule $s^{i}_{tk}:M^{2}\to K$ specifying, respectively, $i$’s effort and technology choice for each possible pair of models held by the two players at the beginning of $t$. A strategy profile $s^{i}=(s^{i}_{1},s^{i}_{2})$ consists of a stage strategy for each $t$. We use the following equilibrium notion.202020 To simplify exposition the analysis focuses on pure strategies, but allowing for mixed strategies does not change our results. Equilibrium existence never requires randomization in the second period. Randomization is never required in equilibrium in the first period if Ann and Bob start the game with the same model. At this level of generality, in specific instances, mixing in the first period might be required for equilibrium existence when players start the game with different models. In those cases, all our propositions still hold, and the inequalities of lemma 4.2 hold for each element in the supports of Ann and Bob’s mixed strategies.
@@
so that, for instance, $\hat{Y}(m^{A},m^{B},Q)>\hat{Y}(m^{A},m^{A},Q)$ means that the expected output of the “most productive” equilibrium is higher when Ann and Bob disagree than when they share Ann’s model. [^21] While the focus on the most productive equilibria provides a simple measure of team performance when there are multiple equilibria, the results of the following sections hold for every equilibrium. Note that, for each $m,m^{\prime}\in M$, $\hat{Y}(m^{\prime},m,Q)=\hat{Y}(m,m^{\prime},Q)$, by symmetry.